1. Discuss the effect of temperature on the spontaneity of reactions with the following values fore ΔH° and ΔS°.
  2. ΔH° = 128 kJ;  ΔS°=  89.5 J/K
  3. ΔH° = -20.4 kJ;  ΔS°=  -156.3 J/K
  4. ΔH° = -127 kJ;  ΔS°=  43.2 J/K
1
Expert's answer
2021-01-13T13:52:29-0500

2.

Enthalpy change = 128 kj = positive value .

Entropy Change =89.5 j /k.mol = Positive Value


So , Gibbs free energy will be " negative "

At " high Temperatures " .

So , Reaction will be Spontaneous at " high Temperature."


3.

Enthalpy change = - 20.4 kj = negative value .

Entropy Change = -156.3 j /k.mol = negative Value.



So , Gibbs free energy will be " negative "

At " Low Temperatures " .

So , Reaction will be Spontaneous at " Low Temperature."

4.



Enthalpy change = -127 kj = negative value .

Entropy Change =43.2 j /k.mol = Positive Value


So , Gibbs free energy will be " negative "

At " All Temperatures " .

So , Reaction will be Spontaneous at " All Temperature." That is Always .